# Sparse Model Soups: A Recipe for Improved Pruning via Model Averaging

### v1

@@ 3.2 Examining Sparse Model Merging
Having demonstrated the advantages of SMS, we now delve deeper into the underlying factors contributing to its success and the cases where it fails.
Exploring parameters beyond random seeds.
In the previous section, we focused on varying the random seed for simplicity. Figure 3 presents a scatter plot for One Shot IMP (70% sparsity) of ResNet-50 trained on ImageNet, comparing the effects of varying the random seed, weight decay strength, retraining duration, and initial learning rate of a linearly decaying schedule. We create eight variations for each setting, selecting equally spaced values around the defaults (details in the appendix). Parameter averages are constructed from two-element pairs of models in the uniform soup setting. The plot displays the test accuracy of the averaged model versus the maximal test accuracy of each pair. In this setting, most averaged models show a net improvement over their individual components, indicating that varying different hyperparameters in the retraining phase of One Shot IMP produces models within the same loss basin. Comparing the different hyperparameters, the most substantial improvement is achieved by varying the random seed, which we attribute to the difficulty of controlling the impact of other parameters: the effect of a different random seed is uncontrollable, while more or less aggressive retraining schedules have more pronounced, specific impacts.
@@
Merging IMP models in later phases.
Throughout our experiments, we observed that in later phases of IMP, assuming regular IMP and not starting from an averaged model as in SMS, sparse connectivities begin to diverge, leading to reduced sparsity when averaging. Specifically, while in the first phase two splits of the models converge to the same loss basin, the next pruning might push them into different subspaces, motivating the combination of individual models into a single one at the end of each phase to ensure that effectively averaging them remains possible. We also experimented with weight alignment, as notably proposed by Ainsworth et al. (2023) and Singh and Jaggi (2020) (see Section 4 for a more detailed account), hoping to permute models at later phases to the same linear subspace. Although we managed to reduce the decrease in sparsity, we were unable to fully recover the original sparsity. This approach can only work if different IMP runs share the same distribution of sparsity among layers (i.e., a specific layer in model 1 must have the same sparsity as the same layer in model 2), which is in general not the case.
@@
Figure 4: WideResNet-20 on CIFAR-100: (a) Accuracy difference between the soup and best performing model after One Shot pruning and retraining. The lines for UniformSoup and GreedySoup show the envelope considering all retraining schedules and durations. (b) Difference between the soup and prolonged IMP version after One Shot pruning to 90% and 98% sparsity. The x-axis shows the retraining length of the $m=3$ models in the soup, while the prolonged IMP version is retrained three times as long. Results are averaged over multiple random seeds with min-max bands indicated.
Exploring limitations: Instability to randomness and recovering it.
@@
Frankle et al. (2020) demonstrate that networks become stable to randomness after a certain amount of pretraining. They specifically analyze the training amount required before splitting up a network into two copies, further trained with different random seeds, such that the final models reside within a linearly connected basin. In that vein, several works (Frankle et al., 2020; Evci et al., 2022; Paul et al., 2023) study the stability of IMP with weight rewinding (IMP-WR) in the context of the Lottery Ticket Hypothesis (Frankle and Carbin, 2018). Evci et al. (2022) demonstrate that trained lottery tickets converge to the same basin as IMP-WR solutions, while Paul et al. (2023) establish that successive IMP-WR solutions with varying sparsity levels are linearly mode connected, ensuring that the error remains relatively stable along the linear interpolation between two adjacent solutions. In contrast, our focus lies on retraining networks without rewinding, where the network undergoes pruning-induced perturbations but is not rewound to a previous state. Given the instability analysis of Frankle et al. (2020), it seems plausible that different splits of a pruned network converge to the same basin when the sparsity is low and the learning rate is not drastically increased. Likewise, we anticipate that high pruning levels might reduce stability to randomness.
3(a) compares the test accuracy difference between a soup of $m=5$ models and the best candidate test accuracy. We vary retraining length (5, 20, and 50 epochs) and retraining schedule (LLR and ALLR (Zimmer et al., 2023)), plotting the top-performing configuration for each sparsity. Both UniformSoup and GreedySoup improve over individual models by up to 2%. However, in the high sparsity regime, the soup benefit starts to diminish, even with extended retraining time. Interestingly, UniformSoup collapses, while GreedySoup maintains performance no worse than the best individual model. This indicates that beyond a certain sparsification level, stability to randomness decreases. At least one of the five models diverges from the others, and merging all five leads to reduced performance. Conversely, GreedySoup selects mergeable candidates, possibly relying on just a single model.
3(b) compares a sparse model soup of $m=3$ against a prolonged IMP version (IMP+), retrained three times as long as indicated on the x-axis. We vary the retraining schedule (LLR and ALLR) and the merging method (UniformSoup and GreedySoup), plotting the top-performing configuration. The key observation is that for relatively low sparsity and thus instability (red line), it is more effective to train $m$ models in parallel and average them than to train a single model $m$-times as long, even with short retraining time. For high sparsities like 98% (green line), averaging $m$ models retrained for a short time is not as good as training a single one for longer. However, there is a break-even point at around twelve epochs, where the benefit of retraining a single model longer vanishes, and the $m$ models are trained enough to be mergeable. In iterative pruning, SMS is expected to need fewer retraining epochs per phase to gain model soup benefits due to the gradual increase in sparsity.
Pruning has previously been found to disproportionately affect different subgroups of the data, amplifying unfairness (Hooker et al., 2019, 2020; Paganini, 2020). In the appendix, we further study the capabilities of SMS to mitigate such pruning-induced unfairness. Our findings indicate that SMS has a less severe negative impact on individual subgroups than IMP.
@@ 3.3 Improving Pruning during Training algorithms
We extend our findings to magnitude-pruning based methods within the pruning during training domain. These methods, unlike IMP, start with a randomly initialized model and sparsify it during regular training. We focus on GMP (Zhu and Gupta, 2017; Gale et al., 2019), DPF (Lin et al., 2020), and BIMP (Zimmer et al., 2023). GMP employs a pruning schedule to dynamically adjust and update a pruning mask, sparsifying the network throughout training. DPF follows the same schedule but enables error compensation by modifying the update rule to use the pruned model’s gradient while updating the dense parameters. BIMP divides the total training budget into a pretraining phase and multiple IMP cycles for the remaining portion, rendering it closest to IMP. As opposed to others, the three approaches can be easily adapted using SMS. BIMP can incorporate SMS into individual phases as previously discussed. Both GMP and DPF prune at uniformly distributed timesteps throughout training. Consequently, we consider the timeframe between two such steps as a phase, during which we create $m$ copies of the recently pruned model, train them in parallel using different random seeds, and merge them before the next pruning step. Unlike IMP or BIMP, no specific retraining schedule ensures convergence in such a phase; instead, GMP and DPF follow the original learning rate schedule.
@@
Table 2 compares the three methods and their SMS-enhanced versions to other state-of-the-art pruning during training methods, including GSM (Ding et al., 2019), DNW (Wortsman et al., 2019), LC (Carreira-Perpinán and Idelbayev, 2018), and DST (Liu et al., 2020). Ensuring a fair comparison we only consider the dense-to-sparse training paradigm, as opposed to pruning at initialization (Lee et al., 2019; Tanaka et al., 2020) and dynamic sparse training methods (Dettmers and Zettlemoyer, 2019; Evci et al., 2020). We applied LC and GSM to both randomly initialized and pretrained models, selecting the best results for each sparsity, noting the original works only applied these to pretrained models. Further, we experimented with STR (Kusupati et al., 2020), but omitted the results as we were unable to cover the exact sparsity range, being controllable only indirectly through regularization parameters. For a complete overview of the training process and specific hyperparameters, please see the appendix.
@@
We find that incorporating SMS into BIMP, GMP, and DPF consistently improves performance, with BIMP improving the most. We attribute this to the decaying learning rate, which allows for convergence within a phase. Secondly, when compared to other methods, SMS emerges as a simple yet effective way to enhance the competitiveness of magnitude-based pruning methods. However, we note that SMS-enhanced methods increase computational cost by branching into $m$ models between phases, a cost which can be mitigated as individual training can be fully parallelized.
@@
Table 2: ResNet-50 on ImageNet: Comparison of BIMP, GMP and DPF with their SMS-extended variants for goal sparsity levels of 70%, 80%, 90%. Each subcolumn denotes the top-1 accuracy and the theoretical speedup achieved by the method. All results are averaged over multiple seeds and include standard deviations. The best, second best, and third best values are highlighted.
@@ 4 Related Work
We review the related deep learning literature on prediction ensembling, model averaging and connectivity of solutions, focusing especially on studies that intersect with sparsity-related themes. We refer to Hoefler et al. (2021) and Ganaie et al. (2021) for a comprehensive review of sparsification and prediction ensembling, respectively.
Prediction Ensembling.
Lakshminarayanan et al. (2017) employ Deep Ensembles to enhance predictive uncertainty, improving model calibration and generalization to out-of-distribution data. Huang et al. (2017) use Snapshot Ensembling, averaging model parameters at different snapshots during training and leveraging a cyclic learning rate schedule to explore multiple local minima. Similarly, Fast Geometric Ensembling (Garipov et al., 2018) finds an ensemble within a single training run, while Mehrtash et al. (2020) and Chandak et al. (2023) build ensembles by perturbing parameters and leveraging validation data gradients, respectively. In the context of sparsity, Liu et al. (2021) use dynamic sparse training (Mocanu et al., 2018; Dettmers and Zettlemoyer, 2019; Evci et al., 2020) methods to efficiently generate diverse networks for ensembling. Whitaker and Whitley (2022) create diverse prediction ensembles by pruning randomly and retraining $m$ copies of a pretrained network, while Kobayashi et al. (2022) finetune different subnetworks of a pretrained network.
Model Averaging.
@@
Stochastic Weight Averaging (Izmailov et al., 2018) enhances model generalization by averaging parameters across the SGD optimization trajectory. Wortsman et al. (2022a) and Rame et al. (2022) demonstrate the effectiveness of model soups, achieved by averaging multiple finetuned models with varied hyperparameters, leading to better performance and out-of-domain generalization. The approach most similar to our proposal in Section 3.3 is Late-phase learning (von Oswald et al., 2020), which independently trains and then averages specific parameters, albeit not in combination with pruning. Gueta et al. (2023) focus on fine-tuning language models, revealing a clustering-like behavior in weight space, with regions around closely clustered models containing potentially superior models. Croce et al. (2023) explore model soups of adversarially-robust classifiers, while Choshen et al. (2022) enhance base models by merging multiple finetuned ones. Wortsman et al. (2022b) propose robust fine-tuning through averaging zero-shot and fine-tuned models. The very recent and concurrent work of Jolicoeur-Martineau et al. (2023) introduced PopulAtion Parameter Averaging (PAPA), which, similar to SMS, averages independently trained models regularly to ensure that the models do not become to dissimilar for effective averaging. While their approach does not incorporate the pruning of averaged models, we think that their results can be viewed as further support for the approach of SMS.
Combining model averaging with sparsity, Yin et al. (2022b) suggest an extension to Dynamic Sparse Training (DST) methods. During the last 10% of training, they apply a cyclic learning rate and further pruning to explore different sparsification patterns, yielding multiple ‘cheap tickets’. These networks are then integrated into a single model, which, despite densifying upon averaging, is pruned to maintain the initial sparsity level. Yin et al. (2022a) proposes averaging adjacent subnetworks of IMP to create superior lottery tickets, with sparsity reduction counteracted by magnitude pruning. Stripelis et al. (2022) introduce a Federated Learning algorithm, FedSparsify, where a global mask is maintained centrally and updated with local client prunings, with disparities resolved by majority voting. Concurrently and independently from our work, Jaiswal et al. (2023) obtain subsequent pruning masks early by exploring and averaging several mask configurations.
Mode Connectivity.
@@
Neyshabur et al. (2020) demonstrate that training two models from scratch leads to models not linearly connected, while models finetuned from a pretrained model tend to be similar and reside within the same loss basin. Entezari et al. (2022) conjecture that two models trained with different seeds are linear mode connected up to permutations of the network’s neurons and filters. Demonstrating this linear mode connectivity, Ainsworth et al. (2023) propose algorithms to find permutations transforming models into a shared loss basin. Singh and Jaggi (2020) leveraged model fusion for neuron soft-alignment across models and improved convolutional filter pruning by fusing the dense model into the sparse one. In that vein, Benzing et al. (2022) propose a permutation algorithm and surprisingly demonstrate that models also lie in the same loss valley (up to permutation) at time of their initialization. Finally, Jordan et al. (2022) explore ‘variance collapse’ in pre-activations when interpolating deep networks, proposing mitigation strategies. There have been several works (Frankle et al., 2020; Evci et al., 2022; Paul et al., 2023) studying IMP’s stability to randomness, we refer to Section 3.2 for a detailed account.

### v3

@@ 3.2 Examining Sparse Model Merging
Having established the merits of SMS, we now investigate its success and limitations in more detail.
@@
Exploring parameters beyond random seeds. Previously, we focused on varying the random seed for simplicity. Figure 3 presents a scatter plot for One Shot IMP (70% sparsity) of ResNet-50 trained on ImageNet, comparing the effects of varying the random seed, weight decay strength, retraining duration, and initial learning rate of a linearly decaying schedule. Exact hyperparameters are listed in Section B.2. Parameter averages are constructed from two-element pairs of models in the uniform soup setting. The plot displays the test accuracy of the averaged model versus the maximal test accuracy of each pair. In this setting, most averaged models show a net improvement over their individual components, demonstrating that varying different hyperparameters in the retraining phase of One Shot IMP produces models within the same loss basin. Comparing different hyperparameters, the most substantial and consistent improvement comes from varying the random seed. Unlike the random seed, which only introduces variability due to inherent randomness, other parameters such as weight decay have a direct, controllable impact on the results; for instance, a poorly chosen weight decay value could significantly degrade performance.
OOD-robustness and fairness. We also explored if SMS enhances robustness to out-of-distribution data, akin to regular model soups (Wortsman et al., 2022a), using benchmark robustness datasets CIFAR-100-C and ImageNet-C (Hendrycks & Dietterich, 2019) for evaluation. SMS consistently outperformed individual models from IMP and other baselines, displaying better resilience to common corruptions, especially with ImageNet-C, which showed up to a 2.5% increase in OOD-accuracy (see Figure 8, Table 14 in Section B.2.3 for more details). Further, previous research suggests that pruning can exacerbate unfairness across data subgroups (Hooker et al., 2019; 2020; Paganini, 2020). In Section B.2.4, we examine if SMS could alleviate such pruning-induced unfairness, and found SMS to exhibit less severe negative impacts on individual subgroups than IMP.
@@
Figure 4: WideResNet-20 on CIFAR-100: (a) Accuracy difference between the soup ($m=5$) and best averaging candidate after One Shot pruning and retraining for varying sparsity levels. (b) Accuracy difference between the soup ($m=3$) and $\text{IMP}_{3\!\times}$ retrained three times as long as indicated on the x-axis, using One Shot pruning to 90%, 95% and 98% sparsity. Results are averaged over multiple random seeds with min-max bands indicated.
@@
Instability to randomness and recovering it. Neyshabur et al. (2020) demonstrated that during training from scratch, the inherent randomness in batch selection alone suffices to cause divergence between two models to the extent that they are not averageable, even when starting from identical (random) initialization. Such instability to randomness can be mitigated by ensuring sufficient pretraining: Frankle et al. (2020) specifically analyze the amount of training required before splitting a network into two copies further trained with different random seeds, such that the final models reside within a linearly connected basin. In that vein, several works (Frankle et al., 2020; Evci et al., 2022; Paul et al., 2023) study the stability of IMP with weight rewinding (IMP-WR) in the context of the Lottery Ticket Hypothesis (Frankle & Carbin, 2018). In contrast, we explore retraining networks without rewinding. Based on the aforementioned instability analysis, we conjecture that different splits of a pruned network converge to a common basin under low sparsity and moderate learning rates, while high pruning levels may potentially reduce stability to randomness.
3(a) shows the difference in test accuracy between a soup of $m=5$ models and the best candidate at different sparsity levels. Each point corresponds to the best configuration across varying retraining lengths (5, 20, and 50 epochs) and schedules (LLR and ALLR). UniformSoup and GreedySoup enhance accuracy by up to 2% over individual models. Yet, as sparsity increases, this benefit declines, with UniformSoup collapsing in performance. Beyond a certain sparsity, stability to randomness declines, causing model divergence and hindering beneficial model averaging. Unlike UniformSoup, GreedySoup performs at least as well as the best individual model.
Similarly, 3(b) depicts the difference between the model soup ($m=3$) and $\text{IMP}_{3\!\times}$, where the latter is retrained for $m\cdot k$ epochs and the former trains $m$ models for $k$ epochs, with $k$ denoted on the x-axis. Again, we plot the best configuration varying the retraining schedule (LLR, ALLR) and merging method (UniformSoup, GreedySoup). For moderate sparsity (green), averaging $m$ models is more effective than training a single model $m$-times as long, even with brief retraining. At high sparsity (red), short-term retraining and averaging $m$ models underperforms compared to extending a single model’s training. However, a break-even point emerges around 15 epochs, beyond which the benefit of extended single model training diminishes, with the $m$ models sufficiently trained for merging. In iterative pruning, we expect SMS to require fewer retraining epochs per phase, benefiting from the gradual sparsity increment. Despite extensive retraining, $\text{IMP}_{m\!\times}$ fails to match SMS.
Efficiency of SMS. Each IMP-cycle consists of $k$ epochs, while $\text{IMP}_{m\!\times}$ extends this to $m\cdot k$ epochs sequentially. Contrarily, SMS executes each phase with $m$ distinct models, independently trained for $k$ epochs, allowing parallelization that can lower wall-time by a factor of $1/m$ compared to $\text{IMP}_{m\!\times}$. Nevertheless, the overall compute and memory requirements for both $\text{IMP}_{m\!\times}$ and SMS remain at $m\cdot k$ units. Resource-wise, $\text{IMP}_{m\!\times}$ and SMS are hence on par; however, the parallelization in SMS underscores a practical advantage. Further, our data shows SMS often significantly outperforming $\text{IMP}_{m\!\times}$, suggesting comparable accuracy can be achieved with fewer total retraining epochs.
In Appendix C, we conduct ablation studies to individually assess the impact of the retraining schedule and the number of epochs per phase during the execution of SMS.
@@ 3.3 Improving Pruning during Training algorithms
We extend our findings to magnitude-pruning based methods within the pruning during training domain. These methods, unlike IMP, start with a randomly initialized model and sparsify it during regular training. We focus on GMP (Zhu & Gupta, 2017; Gale et al., 2019), DPF (Lin et al., 2020), and BIMP (Zimmer et al., 2023). GMP applies a pruning schedule to iteratively update a pruning mask throughout training. DPF, while following the same schedule, enables error compensation by updating the dense parameters using the pruned model’s gradient. BIMP divides the training budget into a pretraining phase and multiple IMP cycles thereafter, rendering it closest to IMP.
These three approaches can be easily adapted using SMS. BIMP can integrate SMS within individual phases. Both GMP and DPF prune at uniformly distributed timesteps during training. We regard the interval between two such steps as a phase, during which we create $m$ copies of the recently pruned model, train them with different random seeds, and merge them before the next pruning step. Unlike IMP or BIMP, GMP and DPF follow the original learning rate schedule throughout a phase.
@@
Table 2 compares accuracy and sparsity-induced theoretical speedup (Blalock et al., 2020) of the three methods and their SMS-enhanced versions to other state-of-the-art pruning during training methods like GSM (Ding et al., 2019), DNW (Wortsman et al., 2019), LC (Carreira-Perpinán & Idelbayev, 2018), and DST (Liu et al., 2020). For a fair comparison, we solely consider the dense-to-sparse training paradigm, as opposed to pruning at initialization (Lee et al., 2019; Tanaka et al., 2020) or dynamic sparse training (DST, Mocanu et al., 2018; Dettmers & Zettlemoyer, 2019; Evci et al., 2020) methods. We applied LC and GSM to both randomly initialized and pretrained models, selecting the best results for each sparsity, noting the original works only applied these to pretrained models. Further, we experimented with STR (Kusupati et al., 2020), but omitted the results as we were unable to cover the exact sparsity range, being controllable only indirectly through regularization parameters. Detailed training and hyperparameters can be found in the corresponding subsection of Section A.2.
@@
Incorporating SMS into BIMP, GMP, and DPF consistently improves performance, despite their deviation from IMP. BIMP benefits the most, likely due to the decaying learning rate facilitating convergence in each phase. In comparison, SMS stands out as a simple yet effective way to enhance the competitiveness of magnitude-based pruning methods, noting that SMS-enhanced methods increase computational costs by branching into $m$ models, a cost mitigable through parallelizing.
@@
Table 2: ResNet-50 on ImageNet: Comparison of BIMP, GMP and DPF with their SMS-extended variants for goal sparsity levels of 70%, 80% and 90%. Each subcolumn denotes the top-1 accuracy and the theoretical speedup at a given sparsity. All results are averaged over multiple seeds and include standard deviations. The best, second best, and third best values are highlighted.
@@ 4 Related Work
We review the related literature, focusing on sparsity-related studies. We refer to Hoefler et al. (2021) for a comprehensive review of sparsification approaches.
@@
Model Averaging. Stochastic Weight Averaging (Izmailov et al., 2018) averages parameters across the SGD trajectory for improved generalization. Wortsman et al. (2022a) and Rame et al. (2022) demonstrate model soups’ enhanced generalization and OOD-performance by averaging models finetuned with varying hyperparameters. The approach closest to ours in Section 3.3 is Late-phase learning (von Oswald et al., 2020), independently training and averaging specific parameters, albeit without pruning. Gueta et al. (2023) explores fine-tuning in language models, revealing a clustering-like behavior with regions around close models containing potentially superior models. Croce et al. (2023) explore soups of adversarially-robust models, while Choshen et al. (2022) enhance base models by merging multiple finetuned ones. Wortsman et al. (2022b) introduce robust fine-tuning through averaging zero-shot and fine-tuned models. Similar to SMS, concurrent work by Jolicoeur-Martineau et al. (2023) regularly averages independently trained models, although without pruning.
We highlight key distinctions between our work and existing studies that combine sparsity with parameter averaging. Yin et al. (2022b) utilize dynamic sparse training, averaging models within a single run with fixed hyperparameters, in contrast to our prune-after-training method that averages models across multiple runs with diverse hyperparameter settings. Their prune-and-grow approach, exploring different sparsity patterns and requiring re-pruning to maintain sparsity, contrasts with our method which deliberately avoids re-pruning by keeping consistent sparsity patterns. We explicitly demonstrate that this approach significantly improves upon the re-pruning approach (IMP-RePrune), even when using strategies like CIA or CAA that are designed to mitigate the impact of re-pruning (Yin et al., 2022b). Similarly, Yin et al. (2022a) employ IMP with weight rewinding, averaging IMP subnetworks of different prune-retrain-cycles across a single training trajectory, which also requires re-pruning, unlike our approach of averaging parallely trained models. Furthermore, their objective is to generate lottery tickets, as opposed to creating sparse models for inference. In a similar vein, Stripelis et al. (2022) introduce FedSparsify, a Federated Learning algorithm that centrally updates a global mask with local client masks, resolving disparities through majority voting. Furthermore, our work is distinct from that of Jaiswal et al. (2023), who focus on averaging early pruning masks for mask generation, whereas we concentrate on averaging parameters of sparse models.
@@
Mode Connectivity. Neyshabur et al. (2020) demonstrate that models trained from scratch are not linearly connected, while models finetuned from a pretrained model tend to be similar and reside within the same loss basin. Entezari et al. (2022) conjecture different-seed trained models are linear mode connected up to neuron permutations. Partially demonstrating this, Ainsworth et al. (2023) propose permutation algorithms for transforming models into a shared loss basin and Singh & Jaggi (2020) employ model fusion for neuron soft-alignment, further also enhancing filter pruning by fusing dense into sparse models. Similarly, Benzing et al. (2022) introduce a permutation algorithm, demonstrating that models share a loss valley (up to permutation) even at initialization. Jordan et al. (2022) explore ‘variance collapse’ in interpolations of deep networks, proposing mitigation strategies. Several works (Frankle et al., 2020; Evci et al., 2022; Paul et al., 2023) study IMP’s stability to randomness, specifically with weight rewinding (IMP-WR). Evci et al. (2022) demonstrate that trained lottery tickets and IMP-WR solutions converge to identical basins, while Paul et al. (2023) find successive IMP-WR solutions at varied sparsity are linearly mode connected, maintaining loss stability along the linear interpolation between adjacent solutions.
Prediction Ensembling. A range of studies focus on prediction ensembling, where outputs of multiple models are averaged (Lakshminarayanan et al., 2017; Huang et al., 2017; Garipov et al., 2018; Mehrtash et al., 2020; Chandak et al., 2023). In the sparsity context, Liu et al. (2021) leverage DST for efficient generation of diverse ensemble candidates. Whitaker & Whitley (2022) form ensembles by randomly pruning and retraining model copies, while Kobayashi et al. (2022) finetune subnetworks of a pretrained model. We refer to Ganaie et al. (2021) for a survey of ensembling.